ROTATIONAL STRUCTURE IN THE B̃-X̃ ORIGIN BAND OF THE METHYL RADICAL

\underline{JEFFREY A. GRAY}, Department of Chemistry, Ohio Northern University, Ada, OH 45810; THOMAS B. SETTERSTEN and ROGER L. FARROW, Combustion Research Facility, Sandia National Labs, Livermore, CA 94551.

The predissociated B̃ ^{2}A′_{1} state of CH_{3} has a sub-picosecond lifetime,^{a} which causes all absorption bands in the B̃ ^{2}A′_{1} ← X̃ ^{2}A′′_{2} system to be unstructured, and molecular constants for the B̃ state must be inferred from analysis of partly resolved lines in the corresponding bands of CD_{3}.^{b} We observe isolated rotational transitions in the B̃ ← X̃ origin band of CH_{3} using two-color resonant four-wave mixing (TC-RFWM) spectroscopy with an IR laser that excites known lines in the ν_{3} fundamental band^{c} to label specific N, K levels of the X̃ state. Analysis of TC-RFWM spectra involves fitting complex Lorentzian lineshapes and yields an updated value of 46, 241±5 cm^{−1} for T_{0} of the B̃ state.

a S. G. Westre, P. B. Kelly, Y. P. Zhang, and L. D. Ziegler, J. Chem. Phys. 94, 270-276 (1991). b G. Herzberg, Proc. R. Soc. London Ser. A 262, 291-317 (1961). c T. Amano, P. F. Bernath, C. Yamada, Y. Endo, and E. Hirota, J. Chem. Phys. 77, 5284-5287 (1982).